# On the Vulnerability of Concept Erasure in Diffusion Models

### v2

@@ 3.2 RECORD
The average gradient $\bar{g}=\frac{1}{J}\sum_{j=1}^{J}g_{j}$ serves as a linear approximation to $\hat{L}$ with respect to the entire prompt embedding space. By multiplying the embedding table $E$ with $\bar{g}$, we efficiently score all possible tokens and select the top $V$ candidates for exact evaluation. During the evaluation and subsequent update of $y_{s}$, we employ a greedy strategy: A candidate token $\hat{c}^{*}$ is only accepted if it improves the loss, i.e. when $\hat{L}(y(\hat{v}^{*},s),\mathcal{R})<\hat{L}(y,\mathcal{R})$. This update process iterates for all positions in the permutation and repeats for $N$ passes, progressively enhancing the token sequence over time. Since each accepted token replacement strictly decreases the loss and the number of possible token sequences is finite, the algorithm is guaranteed to converge to a tangential coordinate-wise local minimum. A pseudocode can be found in Algorithm 1.
@@
Although the RECORD algorithm as stated above is tailored for attacking denoiser-based erasure methods, it can be easily adapted to text-encoder-based erasure methods by replacing $\epsilon_{\theta^{\prime}}\left(z_{t},t,\mathcal{T}(y)\right)$ with $\epsilon_{\theta}\left(z_{t},t,\mathcal{T}^{\prime}(y)\right)$ in equation (2), where $\mathcal{T}^{\prime}$ is the unlearned text encoder.
@@
$\left\lfloor\begin{array}[]{llr}\text{Sample }J\text{ random tokens }\{v_{j}\}&&\\ \text{Compute gradients }\bar{g}=\frac{1}{J}\sum_{j=1}^{J}\nabla{v_{j}}\hat{L}(y({v_{j}},s),\mathcal{Z}(n,s))\\ \text{Score all tokens:}\text{scores}\leftarrow E\,\bar{g}\\ \text{Select top }K\text{ tokens }\mathcal{V}\text{ based on scores}\\ \end{array}\right.$
@@
        if $\hat{L}\big{(}y(\hat{v}^{*},s),\ \mathcal{R}\big{)}<\hat{L}(y,\ \mathcal{R})$:
@@ 4 Experiments
We extensively compare RECORD against the two current state-of-the-art methods in the literature, P4D [3] and UD [47], on text-to-image diffusion models unlearned with both denoiser-based (ESD [10], ED [43], SH [42], FMN [45], CA [20], SPM [25], SalUn [8], UCE [11], and RECE [12]) and text-encoder-based (AdvUnlearn [46]) concept erasure methods. Although there are other black-box methods that suppress generation of harmful content without post-hoc training the denoiser or the text encoder [44, 32, 23, 15], such methods target a fundamentally different adversary, e.g. one with only API access, and cannot exploit internal gradients or weights of the unlearned model. Including them here would not yield an apple-to-apple evaluation of concept erasure under full model access. Unless stated otherwise, all experiments reported in this section aim to find non-seed-specific adversarial prompts, instead of the seed-specific setting commonly investigated in prior literature. The weights of the unlearned models in this section are obtained from [46, 12], which uses Stable Diffusion 1.4 [30] as the base model. The erased concepts include art style (van Gogh), objects (church, garbage truck, parachute) and nudity.
4.1 Evaluation metrics
Most text-to-image diffusion models can generate a far broader range of objects and styles than any single image classifier is capable of classifying. Prior work therefore evaluates concept erasure and restoration methods by ensembling multiple classifiers, each with its own architecture, training data, and preprocessing method, introducing potential inconsistencies. To address this issue, as well as to improve reproducibility and ease replication, we adopt a single, unified zero-shot diffusion classifier (Stable Diffusion v2.1) [22, 4] to classify the generated images. The classification results are obtained by computing
@@ 4.1 Evaluation metrics
where the timestep $t$ is uniformly sampled from $U(0,T)$, $Y=\{y_{1},y_{2},\dots,y_{n}\}$ is a set of $n$ classification classes. The expectation is computed over 10 samples, which in our experience is sufficient to provide accurate classification results. For art style and object attacks, we build sets of 50 classes using prompt templates ‘a painting in the style of {artist_name}’ and ‘a photorealistic image of {object}’, where the artist names are randomly chosen from a list of famous painters, e.g. Leonardo Da Vinci, and object names are from the classification classes of YOLOv3 [29]. For nudity attacks, a set of 4 classes are built with the same template as the object attacks. For all attacks, we also additionally add one empty class, ‘ ’, which helps the classifier capture images that fall significantly outside the distributions of the specified classes. All results presented in this section are computed on 500 images generated by the corresponding erased models and restoration attacks. We report the Attack Success Rate (ASR) in percentage, calculated by dividing the number of images classified as the target (erased) class by the total number of generated images. We attach the classification accuracy of the zero-shot diffusion classifier on images generated by the baseline model in Table 2, which serves as references for the ASRs of the corresponding concepts.
@@
Table 2: The accuracy of Stable Diffusion 2.1 in zero-shot classifying images generated by Stable Diffusion 1.4
4.2 Prompt Attacks
In this experiment, each restoration method (P4D[3], UD[47], RECORD) aims to find 64-token-long non-seed-specific adversarial prompts starting from a randomly initialized prompt sequence, except UD. Since UD optimizes on a token distribution, we follow their original initialization strategy by setting the first few tokens to be the target prompt, and initializing the rest of the tokens from a uniform distribution for all tokens. Without this type of initialization UD does not achieve any significant results. Each restoration methods is evaluated by first creating 50 adversarial prompts. Each prompt is then used to create 10 images, totaling 500 images per method to be used for ASR calculations. For RECORD, we use $N=20$ passes through the token list, a batch size of 1 image each, and $J=64$ samples for the candidate selection. The chosen candidate set has size $K=64$. These numbers are chosen to make the best use the available VRAM of an 80GB H100 GPU. We include results on how runtime scales with number of tokens (Table 3). The runtime of RECORD is competitive despite its parameters are not tuned to minimize runtime. UD has a shorter runtime due to its use of loss (3). This loss can also be used by RECORD, which is a tradeoff between marginal attack performance gain and runtime, as investigated in Appendix A.
Number of tokens Restoration Method Runtime$/s$
P4D UD RECORD
8 $2059\pm 49$ $305\pm 7$ $\textbf{235}\pm 16$
16 $2050\pm 34$ $\textbf{429}\pm 7$ $464\pm 18$
32 $2070\pm 22$ $\textbf{638}\pm 39$ $929\pm 16$
64 $2079\pm 22$ $\textbf{1003}\pm 50$ $1859\pm 15$
Table 3: The runtime comparison is computed over 10 runs. Both $K$ and $J$ are set to the number of tokens.
@@ 64 $2079\pm 22$ $\textbf{1003}\pm 50$ $1859\pm 15$
Table 4: Attack success rate (%) for models with the concepts (van Gogh style, Church, Garbage Truck, Parachute, Nudity) erased, compared between the restoration methods P4D [3], UD [47], and RECORD (ours). The best and second-best values are marked in red and blue, respectively.
@@
The results from Table 4 show that P4D and UD struggle to find non-seed-specific adversarial prompts, with some performing even worse than the no attack baseline, underscoring their ineffectiveness when the seed is not fixed during optimization. On the other hand, RECORD consistently outperforms P4D and UD by up to 17.8 times (see the AdvUnlearn-Parachute result), except for a few minor exceptions. In particular, AdvUnlearn is quite resilient against P4D and UD with single digit ASR on most concepts, while RECORD is able to achieve an ASR of at least $33\%$ for all concepts. SH models are very robust against all concept restoration attacks, but this comes at a significant cost of the generation quality of other non-erased concepts, which has been investigated and discussed in detail by [46].
Example images on van Gogh style can be found in Table 1, as well as in Appendix B for other erased concepts. We also provide an ablation study on the number of tokens $K$ in Appendix C, where we demonstrate RECORD maintains excellent performance even with only 16 tokens (Tables 11,12).
Figure 1: Behavior of the prompt embeddings during prompt embedding attacks. a) Isomap projection of the optimization trajectories of the prompt embeddings during embedding attacks. The rest of the figures are cosine similarity histograms between b) the initial and target embeddings; c) the optimized and target embeddings; d) the optimized and initial embeddings. Target prompt embedding is “a painting in the style of van Gogh” encoded by the non-unlearned text encoder from the baseline model. All data points shown here can successfully generate the erased concept from the unlearned models in the non-seed-specific setting.
4.3 Prompt embedding manifolds
We carefully assess the vulnerability of the concept erasure methods against small perturbations by examining the optimization behaviors with different initializations for prompt embedding attacks. We present a 2D isomap [40] projection (Figure 1a) to visualize the optimization trajectories of prompt embeddings and histograms of the cosine similarity between initial, optimized, and target embeddings (Figure 1b-d). The use of cosine similarity is motivated by the observation that text embeddings typically lie on a hypersphere. We note that our observations are consistent across all denoiser-based concept erasure methods examined, but for clarity we only show ESD (denoiser based) and AdvUnlearn (text encoder based) for comparison. We consider four different initialization schemes: In the first two schemes, initial prompts are sampled “close” to the target prompt of the erased concept by padding the target prompt with random characters ($\pm$ characters) or by padding and replacing random tokens ($\pm$ tokens) of random lengths. On the other hand, initial prompts are sampled “far” from the target prompt by uniformly sampling random characters or tokens.
Embeddings initialized either close or far from the target prompt can successfully restore the erased concept with short optimization trajectories (Figure 1a and b). In particular, prompts initialized in the “far” region do not approach the cluster that contains the target embedding projection. This suggests that, in any region of the prompt embedding space, there almost always exists a nearby prompt embedding capable of restoring the erased concept. In other words, unlearned models are generally vulnerable to small perturbations to the embedding of any arbitrary prompt. Furthermore, the comparison between the initial and optimized text embeddings (Figure 1c and d) reveals that for denoiser-based erasure methods (ESD), embeddings initialized close to the target prompts tend to diverge from the target prompt embedding, whereas, in the case of text-encoder-based erasure methods (AdvUnlearn), embeddings initialized far from the target prompts tend to converge to it. This suggests that denoiser based erasure methods are mainly effective in suppressing the generation of the erased concept at prompts that are embedded close to the target. For AdvUnlearn initial embeddings (Figure 1a), we use the unlearned text encoder, which leads to radically different (from ESD) initialisations for the target prompt and those perturbed on a token level. However, when appending random characters, or initializing with random characters, AdvUnlearn encodes like the baseline encoder. Despite this, both techniques are as vulnerable to small perturbations (Table 5 in appendix A).
Figure 2: Cosine similarities of trajectories of U-Net bottleneck features between unlearned and baseline models during inference, using RECORD-restored (a-b) and non-perturbed (c-d) target prompts. These are compared with the black-colored baseline inference trajectories of semantically related prompts. We borrow the nomenclature from literature [37] and use "Mid" and "Down" to represent the bottleneck and final downsampling layers of the U-Net, respectively. In this particular case, we use a target prompt "a painting in the style of van Gogh", a similar prompt "a painting of starry night by van Gogh", and an unrelated prompt "a photorealistic image of a sports car" as reference trajectories of semantically similar and unrelated text prompts. All generations use the same
pseudo-random number
seed and thus the same latent initialization, unless noted otherwise. For (a-d), only the prompts of the non-baseline models are restored by RECORD, and the trajectory distances with the adversarial prompts are averaged over 20 successful restoration attempts.
4.4 Semantic Latents
Motivated by the prior literature [37, 26], we also compared the bottleneck and final downsampling features of the U-Net denoiser in the unlearned and baseline models to provide a different perspective to the vulnerability of the unlearned models. We compare trajectories of these features during inference with both the target and restored prompts, and present the cosine similarity between the trajectory of the baseline model from the target prompt and the trajectories of the unlearned / baseline models from the restored / target prompts in Figure 2.
Unlearned models with successfully restored concepts have trajectories as far from the baseline-target trajectory as trajectories generated by unrelated prompts in the baseline model (Figure 2a-b), whereas unlearned models with the non-perturbed target prompt produce trajectories much closer to the baseline model with the same target prompt (Figure 2c-d), despite not rendering van-Gogh-related images. These findings suggest that there almost always exists some inference trajectories far away from that of the baseline, which an unlearned model can take to generate the unlearned concept. Consequently, the denoiser-based unlearning methods do not remove information of the concept from the denoiser, but again merely suppress their generation from prompts close to the target prompt.
@@ 5 Conclusions
In this study, we introduced RECORD, a novel white-box concept restoration algorithm designed for restoring erased concepts from unlearned text-to-image diffusion models by adversarially perturbing the input prompts. To the best of our knowledge, RECORD is the first algorithm capable of consistently carrying out non-seed-specific concept restoration attacks, outperforming existing restoration methods by up to 17.8 times. Furthermore, our investigation into existing concept erasure techniques revealed that adversarial prompts are pervasive throughout the prompt embedding space. These insights not only underscore a significant vulnerability inherent in current erasure approaches but also pave the way for future research, inspiring the development of new concept erasure and restoration algorithms that can more effectively mitigate or exploit these vulnerabilities.
Acknowledgment
LB, AR, and KZ were supported by the EPSRC Centre for Doctoral Training in Mathematical Modelling, Analysis and Computation (MAC-MIGS) funded by the UK Engineering and Physical Sciences Research Council (grant EP/S023291/1), Heriot-Watt University and the University of Edinburgh. This work was supported by the Edinburgh International Data Facility (EIDF) and the Data-Driven Innovation Programme at the University of Edinburgh. We thank Christos Christodoulopoulos for his useful discussions, and for his input to the Hackathon Workshop on Generative Modeling.
Appendix A Choice of loss function
We consider two loss functions from literature [27, 47, 3], that are commonly used for identifying adversarial prompts for restoring concepts from unlearned models:

### v3

@@ 3.3 Token-Level Attacks: New Method
The average gradient $\bar{g}=\frac{1}{J}\sum_{j=1}^{J}g_{j}$ serves as a linear approximation to $\hat{L}$ with respect to the entire prompt embedding space. By multiplying the embedding table $E$ with $\bar{g}$, we efficiently score all possible tokens and select the top $K$ candidates for exact evaluation. This effectively alleviates the intractability introduced by the large vocabulary in the naive approach. During the evaluation and subsequent update of $y_{s}$, we employ a greedy strategy: a candidate token $\hat{v}^{*}$ is only accepted if it improves the loss, i.e. when $\hat{L}(\mathcal{T}(y(\hat{v}^{*},s)),\mathcal{R})<\hat{L}(\mathcal{T}(y),\mathcal{R})$, where $\mathcal{R}$ is the reference set. This update process iterates through all positions in the permutation and repeats for $N$ passes, progressively enhancing the token sequence over time. Since each accepted token replacement strictly decreases the loss and the number of possible token sequences is finite, the algorithm is guaranteed to converge to a tangential coordinate-wise local minimum. A pseudocode can be found in Algorithm 1.
@@
Although the RECORD algorithm as described above is tailored for attacking denoiser-based erasure methods, it can be easily adapted to text-encoder-based erasure methods by replacing $\mathcal{T}$ with $\mathcal{T}^{\prime}$ in $\hat{L}$ when encoding $y$, where $\mathcal{T}^{\prime}$ is the unlearned text encoder.
@@
[…]
@@
   if $\hat{L}\big(\mathcal{T}(y(\hat{v}^{*},s)),\ \mathcal{R}\big)<\hat{L}(\mathcal{T}(y),\ \mathcal{R})$:
@@ 4 Experiments
We designed our experiments to address the following questions:
  1. 1.
    Does the proposed RECORD algorithm outperform current SOTA concept restoration methods?
  2. 2.
    Are certain concept erasure methods more robust?
  3. 3.
    How important are the different RECORD hyperparameters?
We extensively compare RECORD against the two current state-of-the-art concept restoration methods in the literature, P4D Chin et al. (2023) and UD Zhang et al. (2023), on text-to-image diffusion models unlearned with both denoiser-based (ESD Gandikota et al. (2023a), ED Wu et al. (2024), SH Wu & Harandi (2024), FMN Zhang et al. (2024a), CA Kumari et al. (2023), SPM Lyu et al. (2024), SalUn Fan et al. (2023), UCE Gandikota et al. (2023b), and RECE Gong et al. (2024)) and text-encoder-based (AdvUnlearn Zhang et al. (2024b)) concept erasure methods. Although there are other black-box methods that suppress generation of harmful content without post-hoc training the denoiser or the text encoder Yoon et al. (2024); Schramowski et al. (2023); Li et al. (2024); Jain et al. (2024), such methods target a fundamentally different adversary, e.g. one with only API access, and cannot exploit internal gradients or weights of the unlearned model. Including them here would not yield an apple-to-apple evaluation of concept erasure under full model access. We use open-sourced unlearned model weights from Zhang et al. (2024b); Gong et al. (2024), which uses Stable Diffusion 1.4 (SD1.4) Rombach et al. (2022) as the base model. The erased concepts include art style (van Gogh), objects (church, garbage truck, parachute) and nudity.
4.1 Evaluation Metric
Most text-to-image diffusion models can generate a far broader range of objects and styles than any single image classifier is capable of classifying. Prior work therefore evaluates concept erasure and restoration methods by ensembling multiple classifiers, each with its own architecture, training data, and preprocessing method, introducing potential inconsistencies. To address this issue, as well as to improve reproducibility and ease replication, we adopt a single, unified zero-shot diffusion classifier (Stable Diffusion v2.1) Li et al. (2023); Clark & Jaini (2023). The classification results are obtained by computing
@@ 4.1 Evaluation Metric
where the timestep $t$ is uniformly sampled from $U(0,T)$, $Y=\{y_{1},y_{2},\dots,y_{n}\}$ is a set of $n$ classification classes. The expectation is computed over 10 samples, which in our experience is sufficient to provide accurate classification results. For art style and object attacks, we build sets of 50 classes using prompt templates ‘a painting in the style of {artist_name}’ and ‘a photorealistic image of {object}’, where the artist names are randomly chosen from a list of famous painters, e.g. Leonardo Da Vinci, and object names are from the classification classes of YOLOv3 Redmon & Farhadi (2018). For nudity attacks, a set of 4 classes are built with the same template as the object attacks. For all attacks, we also additionally add one empty class, ‘ ’, which helps the classifier capture images that fall significantly outside the distributions of the specified classes. All results presented in this section are computed on 500 images generated by the corresponding erased models and restoration attacks. We report the Attack Success Rate (ASR) in percentage, calculated by dividing the number of images classified as the target (erased) class by the total number of generated images. We attach the classification accuracy of the zero-shot diffusion classifier on images generated by the baseline model in Table 2, which serves as references for the ASRs of the corresponding concepts.
@@
Table 2: The classification accuracy of Stable Diffusion 2.1 as a zero-shot image classifier
4.2 Results
In this experiment, each restoration method (P4D Chin et al. (2023), UD Zhang et al. (2023), RECORD) aims to find 64-token-long seed-agnostic adversarial prompts starting from a randomly initialized prompt sequence, except UD. Since UD optimizes on a token distribution, we follow their original initialization strategy by setting the first few tokens to be the target prompt, and initializing the rest of the tokens from a uniform distribution for all tokens. Without this type of initialization UD does not achieve any significant results. Each restoration method is evaluated by identifying 50 adversarial prompts on an H100 GPU and using which to generate 500 images per method for ASR calculations. For RECORD, we use $N=20$ passes through the token list, a batch size of 1 image each, and $J=64$ samples for the candidate selection. The chosen candidate set has size $K=64$. Example images can be found in Table 1 and Appendix J.
@@ 4.2 Results
Table 3: Attack success rate (%) for models erased on different concepts (van Gogh style, Church, Garbage Truck, Parachute, Nudity), compared with different restoration methods P4D Chin et al. (2023), UD Zhang et al. (2023), and RECORD. The best and second-best values are marked in red and gray, respectively.
@@
RECORD consistently outperforms P4D and UD (Table 3) by up to 17.8 times (see the AdvUnlearn-Parachute cells), except for a few minor exceptions. In particular, AdvUnlearn is quite resilient against P4D and UD with single digit ASR on most concepts, while RECORD is able to achieve an ASR of at least $33\%$ for all concepts. Additionally, different erasure methods seem to have different level of robustness against adversarial attacks on different concepts. For example, ED Wu et al. (2024) is more robust in erasing nudity-related concepts, but not on objects such as church. Similar observation can also be observed in Appendix F. On the other hand, SH models are very robust against all concept restoration attacks, but this comes at a significant cost of the generation quality of other non-erased concepts, which has been discussed in detail by Zhang et al. (2024b). We additionally conduct fixed-seed ablation study and demonstrate RECORD still perform well in Appendix E.
5 Discussions
To validate the design of RECORD, we conducted several ablation studies. First, we addressed the ongoing debate in the literature regarding the optimal loss function for concept restoration Pham et al. (2023); Zhang et al. (2023); Chin et al. (2023). Our findings in Appendix B show that the loss function in Equation [1], which uses the original model’s noise prediction as a target, marginally outperforms alternatives that rely on Gaussian noise. This is because the original model’s predictions act as a more informative surrogate, justifying its slight increase in computational overhead.
The runtime of RECORD is very competitive with existing restoration methods (Figure 3), and its hyperparameters $S,N,J,K$ provide flexible control over the compute-performance trade-off (Appendices C,D). For example, our ablation studies in Appendix D demonstrate that a significant 60% acceleration can be achieved by lowering the number of gradient tokens $J$ with only a marginal loss in performance. This allows RECORD’s runtime to surpass that of P4D and UD while maintaining its superior ASRs. Additionally, our studies on prompt length $S$ and the number of passes $N$ conclude that the best strategy for maximizing ASR for a given compute budget is to choose higher $S$ and lower $N$, as the breath of the search space is more significant than the depth. These results can be found in Appendix C.
Figure 3: The mean runtime and the standard deviation (annotated as error bars) of different restoration methods, computed over 10 runs and at sequence lengths $S=8,16,32,64$. We note it is possible to achieve substantial acceleration by lowering gradient token number $J$, with only marginal performance loss, as discussed in Appendix D.
We also tested the transferability of successful adversarial prompts from SD1.4 to larger unlearned models like SDXL Podell et al. (2024) and FLUX Labs et al. (2025) in Appendix F. Our results show that adversarial prompts identified by RECORD are generally more transferable than those from P4D and UD. This finding has significant black-box implications, as an adversary could use a prompt optimized on an open-sourced model to attack a different black box model without requiring excess to the model weights. This superior level of transferability allows RECORD to mitigate the limitations of the white-box assumption commonly used by the existing concept restoration methods.
Lastly, we investigate the scalability of RECORD on a more challenging setting i.e. restoring erased concepts on more sophisticated text-to-image models, such as SDXL and FLUX. In particular, these larger models have a dual-encoder setup for encoding prompts, which differs from the single text encoder used in SD1.4. This introduces additional difficulties in restoration algorithm design and is rarely addressed in the existing concept restoration literature Zhang et al. (2025). We assess five different strategies for adapting RECORD to handle this architectural difference and provide some initial results (Appendix G). Example images by SDXL and FLUX are showcased in Appendix K.
@@ 6 Conclusions
In this study, our investigation into existing concept erasure methods used in text-to-image diffusion models reveal that adversarial prompt embeddings are pervasive throughout the embedding space, which can be exploited by token-level concept restoration methods. We further introduce RECORD, a novel token-level concept restoration algorithm designed for restoring erased concepts by adversarially perturbing the input prompts in a coordinate-descent manner. We conduct extensive experiments and ablation studies demonstrating the consistent superiority of RECORD as well as the effect of its hyperparameters. These results not only underscore significant vulnerability inherent in current erasure approaches, but also pave the way for the future development of erasure and restoration algorithms that can more effectively mitigate or exploit these vulnerabilities.
Appendix A "Vulnerability" Analysis on the Original Model
Figure 4: Behavior of the text embeddings during embedding-level attacks on original model. a) Isomap projection of the optimization trajectories of the prompt embeddings. Dots denote the starting point of the trajectory. The target concept can be generated at the end of each trajectory. b), c), d) are smoothened cosine similarity histograms between the initial, optimized, and target embeddings.
Extending the discussion from Section 3.2, which established the pervasive existence of adversarial embeddings in unlearned models, this section investigates the fundamental origin of this vulnerability. We seek to answer a critical question: are these vulnerabilities an artifact of the concept erasure process, or are they an intrinsic property of the original, pre-unlearned diffusion model?
To explore this, we conduct an experiment on the original, pre-unlearned model using the same embedding-level optimization methodology described in Section 3.2. The objective is to determine if prompt embeddings capable of generating a specific target concept could be discovered starting from a completely irrelevant and semantically distant initial concept.
We use "a painting in the style of van Gogh" as the target prompt, and "a photorealistic image of a sports car" as the semantically distant starting point. We generate 100 paraphrases of the starting prompt and target prompt using Gemini 2.5 and encode them as the embedding initializations and as the reference cluster, respectively.
Our findings reveal that even when initialized with embeddings for a "sports car," the optimization process consistently discovers local embeddings that can generate van Gogh style paintings, far from the region associated with the actual "van Gogh" prompt (Figure 4).
This result provides a crucial insight: the high-dimensional embedding space of the original model is densely populated with regions that can trigger the generation of a given concept. The vulnerability to concept restoration is therefore not primarily induced by the unlearning process but is an characteristic inherited from the pre-unlearned model. Most concept erasure techniques focus on suppressing a concept in the localized region around its explicit text description, leaving these numerous, distant regions of vulnerabilities adversarially exploitable by the restoration attacks analyzed in this paper.
Appendix B Embedding-level Attack Results and Loss Function Comparison
The selection of an optimal loss function for identifying adversarial prompts to restore concepts from unlearned models is a topic of active debate Pham et al. (2023); Zhang et al. (2023); Chin et al. (2023). This ablation study seeks to resolve this ambiguity by empirically comparing two prominent loss functions: